A crane is used to lift some building materials which weigh 112 kg from the ground to the top of a building.





If the building materials are being raised by an electrical winch at a speed of 2.13 m s−1 then at what rate is the winch using electrical power? (in W to 3 s.f)





note: use g = 9.81 m s−2.






1
Expert's answer
2022-04-11T12:04:17-0400

The electrical power

"P=\frac{W}{t}=\frac{mgh}{t}=mgv"

"P=112*9.81*2.13=2340\:\rm W"